In a rectangle, the length is fifthteen less than three times the width. If the perimeter is 34, what is the length?
Answer by mananth(16946)


Perimeter = 34 feet
Length 1 ( 3 x -15 )
Width = 1 x ft
Perimeter =2(L+W)
Perimeter = 2((3x-15)+x)
Perimeter = 2((3x-15+x)
Perimeter = 2(6x-15+2x)
34=16x-30
34-=16x
34+30=8x
64=8x
x=8ft Width
plug x in the expression for length(3x-15)
we get( 24-15)
Length= =9 feet